Minutes of the Trigger Operation Meeting (24th July, 2007)
Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Su Dong, Mark Tibbetts.
- Sat pm: paged for TSF damage. It was of the minor kind, so I instructed the pilot to end the run and start a new one which cleared the problem.
- During the PEP troubles on Monday the timing signal was lost, which required a reboot of the TSF crates. Chris O'Grady took care of it.
- See plots here. First is M turn-on curve for mu pairs from two runs taken at the beginning of July, one with high rate (4.3kHz in red) and one with a low rate (1.4kHz in black).
The MIP energy is 180GeV, and the difference in efficiency at that point is less than 1%.
It was previously observed that EMT efficiency did not drop over time, and here we can see that there is also little change in efficiency with high background.
Interestingly, high background seems to result in an effective shift in energy threshold to higher energy. Further investigation?
The second plot shows the timing for each channel in one mu pair event.
- The 1E trigger line will be included at some point soon. Some discussion is required on what L3 changes are needed - should 1E be included in the physics stream? If so, how will the Bhabha vetoes affect it? It was tentatively decided that 1E should be put through as a separate stream and not included in the L3InPhysics stream.
- It appears that an effect of the EMC L3 filter is cutting out some possibly interesting two-photon physics. There is a requirement on the total invarient mass >1.5GeV which cuts out the gammagamma->f_980 resonance.
A possibility is to produce a special filter for this analysis.
Data Quality and Monitoring
- No problems so far with trigger data.
This page is created by Debbie Bard (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.